The first thing to do were Flats for Ha first for the previous night but also for use tonight. I completed this at 22:15
At 22:31 I performed focus run on Luminance which came in at position 59841 with HFR 4.77.
I started an image run as soon as the Pelican Nebula (IC5067) was above the neighbours house. I have set a run of 10 x 600s and 20 x 300s Ha subs guided with the new PHD2 settings to prove the 12″ now works from 1 night to another. Then we can try the Esprit on the next clear night. The guiding at first looked okay.
The resulting image looked very good too
After about 5 frames I suddenly ran into a problem the guiding looked like this
There were large movements in the Dec and the RA would not return to centre. Soon enough SGPro stopped imaging and tried to settle the guider and then further complained about not being able to settle. So I temporarily changed the Hysterisis from 10 to 15 to see if this would get the star back. It did, well just.
The RA axis returns to the centre after quite some time. Is this possibly seeing related or is there something mechanical amiss with the setup?
I did notice the problem one more and then went off to bed left it running 01:09 leaving the possible bump in the worm to resolve itself.
Addendum – So got up this morning and the scope had continued its travel across the heavens. I looked at SGPro and it finished the sequence without any problem. I then looked at PHD2 and expanded the time to include the maximum span possible as it had been trying to guide after loosing the star (I have check boxed Enable Star Mass Detection in the hope this fixes it) and I noticed a problem, which is the recurring South then North oscillation which I now need to investigate as that is the cause of loosing an image due to the star moving. It looks for all the world to be on the worm gear. I will set about measuring the PEC tonight and see if I can see it. It may of course be from when I adjusted the work due to another problem I had and it may not be quite right. I will also ask on the forum.
M94 and NGC 3395/3396 are the 2 targets for tonight, some luminance on M94 and RGB on NGC 3395/3396 if I get enough time. I always try to open the dome early to give at least 2-4 hours cooling before I use.
When I was about to start with autofocus I tried to recenter back on the target but the mount respond and it transpired that the mount thought it was out of balance. I went to the dome and the mount was beeping proving it was out of balance. So I turned the mount off, manually moved the scopes pack to the park position and then turned the mount back on and all was well.
@19:57 I performed the autofocus for the night on Luminance which scammer in at a position of 75282 on the focuser.
@20:10 I started an imaging run of 24 x NGC 3395/3396 with Luminance filter. Once done I planned on grabbing RGB frames before moving on to M94.
@22:32 I started on the RGB frames for NGC 3395/3396 after refocusing on the Red filter.
@1:40 I slewed to M94 and changed the filter to Luminance. I performed a refocus and shifted from 77895 to 75884 on Red filter by accident. So we (I had Bob on Zoom by this point) refocused on the Luminance and the new focus position was 74884. So the difference is 1000 for Luminance to Red. I also changed the step size for the focuser temperature compensation from 531 to 431 to see if the HFR is more stable.
I noticed tonight that PHD2 lost the Use Direct Guide check mark twice and thus complained about pulse guide not being supported. I had to stop guiding, disconnect the mount in PHD2 and go into the settings, check the Use Direct Guide and reconnect the mount and start guiding again. Something to look into possibly.
@02:07 I went to bed and left the scope gathering another 2 hours of Luminance data on M94.
The following day I took the ZWO ASI1600MM CMOS Camera off the back of the 12″ and cleaned the sensor window. What I found was the dark dust doughnuts disappeared and the rest for the doughnuts were actually on the filters.
Things to still resolve……..
Check out why WSX is loosing connection and shutting the dome
Fix Slew Here and Centre Here in SGPro that does not work